A CAR TRAVELLED 1.98KM. FIND THE NUMBER OF REVOLUTION MADE IF THE RADIUS OF WHEEL IS 105 CM
Answer by Alan3354(69443)

1.98 km = 1980 meters.
----
105 cm = 1.05 meters
Circumference = 2*pi*r = 2.1*pi meters
------------
Revolutions = 1980/(2.1*pi)
===================
PS 2.1 meter diameter is a big wheel.